OpenStudy (hannahdeanm):

Arrange the circles (represented by their equations in general form) in ascending order of their radius lengths. x2 + y2 − 2x + 2y − 1 = 0 x2 + y2 − 4x + 4y − 10 = 0 x2 + y2 − 8x − 6y − 20 = 0 4x2 + 4y2 + 16x + 24y − 40 = 0 5x2 + 5y2 − 20x + 30y + 40 = 0 2x2 + 2y2 − 28x − 32y − 8 = 0 x2 + y2 + 12x − 2y − 9 = 0

OpenStudy (mathmale):

Wow. You're encircled by circles!! Looks like you'll have to go through all of the equations and determine the radii, then rearrange the circles in ascending order by radius length. First problem: x2 + y2 − 2x + 2y − 1 = 0 Note: Please use x^2 (not x2) to denote "the square of x." Rewriting this by grouping the x-terms and the y-terms:\[x^2 -2x +y^2 +2y = 1\]

OpenStudy (mathmale):

We have to "complete the square for both the x^2 -2x and the y^2+2y groups. Know how to do this? Starting with x^2-2x, take half of the coefficinet of x and square it. In other words, take half of -2, which gives us -1, and square -1, to get +1. Add, then subtract, 1 from x^2-2x: x^2-2x+1-1. Perform the same operations on the y terms: y^2+2y+1-1 Rewrite all this as \[x^2-2x+1 + y^2 +2y + 1 -1 -1 =1, \]

OpenStudy (mathmale):

... or ...\[(x-1)^2+(y+1)^2 =3 = (\sqrt{3})^2\]

OpenStudy (mathmale):

The radius of this circle is 2. The center of the circle is at (1, -1). Please try the next equation. If you need more info and/or practice on how to "complete the square," please ask.

OpenStudy (hannahdeanm):

thanks. i understand how to do it now @mathmale

OpenStudy (mathmale):

I'm really happy for you. Good luck!